# Rethinking Weight-Averaged Model-merging

### v3

@@ 6.2.1 Observations over Different Models
As shown in Fig. 3 for the merging of 10 models (the experiments for merging from 2 to 7 models are in the supplementary materials). For abbreviations, “Perf Ave” denotes the average performance of individual models; “deit_tiny” and “vit_tiny” denote deit_tiny_patch16_224 and vit_tiny_patch16_224 models, respectively. Different models show different merging/ensembling performances, but model ensemble on logits generally performs the best across different tested architectures [9, 10, 24, 26, 5, 29]. However, uniform soups perform the worst. We can also discover that greedy soups generally perform better than uniform soups due to their mechanism to greedily select the best performing model to merge. In fact, it guarantees that reasonably promising results can be achieved by only incorporating the first model (if the later-on merged model performance is not ideal). The greedy ensemble on both logits and features produces results that are comparable to greedy soups.
In general, adding more models into the merging process leads to better performance for all methods, including model-merging and ensembling. Interestingly, the ViT model, including 3 ViT-based architectures in the experiments, shows large gaps between uniformly averaging and logits ensembling, where such gaps suffer larger increases by adding more models into. This happens may be caused by the architectural differences between convolutional neural networks and Vision Transformers (ViTs). CNNs are more adapted to the learning of local patterns within different scales of images, making it ideal for merging. On the contrary, ViTs have global attention mechanisms, where the templates represent attention and correlations between tokens, making them sensitive for small changes. ViT models also show similar deterioration on other data (as shown by the DeiT on TinyImageNet and it can be found in the supplementary materials).
@@ 6.2.2 Observations over Different Datasets
Table 1: Performance Comparison for VGG19 with 10 Models Across Different Datasets. We adopt averaged AUROC ( Area Under Curve of Receiver Operating Characteristic) as the evaluation metric (normalized into 0 to 100 for better representation with other datasets) for the ChestXRay [33, 18] dataset; while accuracy (0% to 100%) for the rest of datasets.
@@
For the dataset-wise examination, we perform model-merging and ensembling on 7 prevalent datasets with different data sizes (TinyImageNet and ChestXRay are relatively large, and both with more than 100,000 images in the datasets). From Tab. 1, we can observe that the ensembling still generally performs better than merging. The models on different datasets have different performances due to the dataset characteristics. Similarly to the model-wise comparison, with more models in the merging process (2 to 7 models merging results are in the supplementary materials), the better the performance for both merging and ensembling. The ensembling is more stable than merging in terms of the best and worst performance on different datasets. For example, on the PathMNIST [39, 40], uniform soup does not work very well (32.42% accuracy, which largely behind model ensemble methods) compared to its performance on the Cifar10/Cifar100 or Tiny ImageNet datasets. A similar situation shows on Celeba [17]: the uniformly average model only achieves 53.19%. Nevertheless, the greedy soup models can mitigate this problem (88.48% on PathMNIST and 93.06 on Celeba) as they can ensure a decent result by only considering the first model. These results may be caused by the contradictions between model weights.
6.2.3 Merging Models trained over different Tasks
Datasets Data Size Model2 Uniform Soups Ens Lgt Ens Features
Cifar10 60,000 5.84 49.63 69.93 69.93
PathMNIST 107,180 1.29 24.25 69.88 70.00
DermaMNIST 10,015 5.77 60.37 69.06 69.36
Celeba 202,599 1.03 49.33 69.48 69.61
TinyImageNet 100,000 18.16 61.02 58.40 58.52
ChestXRay 112,120 15.18 61.38 68.18 68.33
We also explore model-merging between different well-trained models on different tasks, as shown in Tab. 2. Model 1 is a VGG19 model trained on the Cifar100 with the classification accuracy of 70.37%, while Model 2 denotes the VGG19 model trained on different datasets, but evaluated on Cifar100 after merging/ensembling. Generally, uniform souping within Tiny ImageNet, where Model 2 was trained on, has the highest accuracy of 61.02% that is better than the ensemble on logits (58.40% accuracy). It may be because the Tiny ImageNet dataset is a general and relatively large dataset, forcing Model 2 to learn a diverse set of features. Interestingly, Cifar10 should be a similar one to Cifar100, but the souping result does not outperform the model ensembling. Maybe because it is too small. In contrast, the ChestXRay dataset should be very different from Cifar100, but it shows a reasonably decent improvement with souping compared to Model 2 along. It illustrates that the dataset size also matters. The souping performance on Celeba (Model 2 is trained on) is not good, may be caused by the Celeba dataset only containing human faces that are very different from Cifar100 (Model 1 is trained on).
@@ 6.3 Model Predictions and Weight Magnitudes
We further visualize the gaps between model-merging and ensembling in Fig. 8(c). When more models are taken into account, the gaps become larger, but there is a peak point for the gap (×$\times 100$ is the peak for Cifar100). The experiments of merging/ensembling 2 to 10 models with different magnitude factors, as well as on different datasets and the gaps between merging and ensembling, are shown in the supplementary materials. A similar phenomenon is shown on different datasets, but with different peaks. Moreover, we conduct the weight magnitude experiment on a relatively large dataset, Tiny ImageNet, in Tab. 3. From the table, we can observe that model-merging can still mitigate the weight magnifying problem to a certain extent. It can also outperform logit ensembling, and cannot outperform feature ensembling. But generally it cannot work very well as on Cifar100. It may be caused by the complexity of the Tiny ImageNet and as we discussed in the Sec. 5.3, the model-merging can work to some extent as it can act similarly as regularization. It can mitigate the weight magnitude increment issues, but with the cost of losing model expressiveness.
Table 3: Model-merging to address the weight magnitude issue for VGG19 with different model numbers on Tiny ImageNet dataset with the magnitude factor ×$\times 100$. This approach works for smaller data (Cifar10/Cifar100) as it has a regularization effect. However, it does not work for larger datasets such as Tiny ImageNet, because it sacrifices the expressiveness of the models.
2 models 59.87 5.60 6.70 6.99 1.59 6.94 6.45 1.52 6.96
3 models 59.85 5.44 7.15 7.05 0.78 7.38 6.78 0.78 7.21
5 models 59.89 5.07 6.68 6.83 0.30 6.95 6.46 0.35 7.21
7 models 59.90 4.80 6.77 6.80 0.24 6.73 6.26 0.21 6.74
10 models 59.90 4.81 6.86 6.92 0.13 6.92 6.22 0.11 6.95
@@ 7 Conclusion
Supplementary Material
8 Experimental Settings
@@
Datasets Sizes Class # Links
Cifar10 60,000 10 https://www.cs.toronto.edu/~kriz/cifar.html
Cifar100 60,000 100 https://www.cs.toronto.edu/~kriz/cifar.html
PathMNIST 107,180 9 https://medmnist.com/
DermaMNIST 10,015 7 https://medmnist.com/
CelebA 202,599 2 https://mmlab.ie.cuhk.edu.hk/projects/CelebA.html
TinyImageNet 100,000 200 https://huggingface.co/datasets/zh-plus/tiny-imagenet
ChestXRay 112,120 14$\times$2 https://www.kaggle.com/paultimothymooney/chest-xray-pneumonia
For the experimental settings of model weight patterns. We train one single linear layer on both Cifar100 [14] and Tiny ImageNet [15] datasets with momentum SGD as the optimizer. The initial learning rater is set to 0.01 and decreases every 20 epochs by 0.1. While visualization, the weights are wrapped back to the image size as described in the main paper (i.e. 100 $\times$ 3 $\times$ 32 $\times$ 32 for Cifar100).
In terms of averaging on weights versus averaging on features, we consider Uniform Soups and Greedy Soups [35] for weight-averaged model-merging, which is the focus of our paper. We design the experiments of this section from two different angles: model-wise examination and dataset-wise examination of the model-merging and model ensemble performances. For the models considered, we have ResNet [9] (ResNet18 to ResNet152), DenseNet [10] (DenseNet121 to DenseNet201), VGG [24] (VGG11 to VGG19, including with and without batch normalization), ViT [26, 5] and DeiT [29]. For the dataset-wise examination, we conduct experiments on Cifar10/Cifar100, PathMNIST/DermaMNIST [39, 40], Celeba [17], Tiny ImageNet and ChestXRay [33, 18]. The information of these datasets can be find in Tab. 4. Specially, ChestXRay-14 is a chest X-ray imaging dataset designed for multi-label classification. It consists of 112,120 frontal chest X-ray images from 30,805 unique patients. The dataset includes annotations for 14 different lung disease categories. We use the official dataset split for training and evaluation. We demonstrate the average Area Under the Receiver Operating Characteristic Curve (AUROC) across 14 labels as the model performance for ChestXRay.
@@
9 Theoretical Proofs
@@
For the Property 1 and Theorem 1, according to [34], we have:
@@
10 More Visualization of Model Weight Patterns
@@
11 More Experimental Results of Averaging on Weights V.S. Averaging on Features
Figure 18: The 2 model merging/ensembling accuracy comparison between different architectures on Cifar100.
Figure 19: The 3 model merging/ensembling accuracy comparison between different architectures on Cifar100.
Figure 20: The 5 model merging/ensembling accuracy comparison between different architectures on Cifar100.
Figure 21: The 7 model merging/ensembling accuracy comparison between different architectures on Cifar100.
The 2 models, 3 models, 5 models and 7 models for model merging/ensembling accuracy comparison between different architectures on Cifar100 are shown via Fig. 18, Fig. 19, Fig. 20 and Fig. 21, respectively. They all show that the models with ensembling on logits generally receive the best results and uniform soups receive the worst results. Besides, model-merging performs much worse results for the ViT models and the gaps between uniform soups and ensemble on logits are enlarged with more models considered in.
@@
Also, we show the dataset-wise exploration of 2 models, 3 models, 5 models and 7 models merging/ensembling in Tab. 5, Tab. 6, Tab. 7 and Tab. 8, respectively. The computation consumptions in Floating Point Operations (FLOPs) and Parameter Numbers are in Tab. 9.
@@
We illustrate the ViT model (DeiT Tiny) on another dataset — Tiny ImageNet in Tab. 10. We can also see the more models are merged, the worse results from uniform soups can get. Also, we can observe the greedy soups, greedy ensemble on logits and greedy ensemble on features are not changed. It demonstrates the model-merging between 2 or more ViT models will not contribute to the performance. Thus, it shows the ViT models are not suitable for model-merging in the tested scenarios.
12 More Experimental Results of Model Magnitude Changes
Figure 23: Accuracy vs. Magnification Factor across Methods with 2 models on Cifar100 dataset.
Figure 24: Accuracy vs. Magnification Factor across Methods with 3 models on Cifar100 dataset.
Figure 25: Accuracy vs. Magnification Factor across Methods with 5 models on Cifar100 dataset.
Figure 26: Accuracy vs. Magnification Factor across Methods with 7 models on Cifar100 dataset.
The 2 models, 3 models, 5 models and 7 models accuracy vs. different magnification factors on Cifar100 are shown via Fig. 23, Fig. 24, Fig. 25 and Fig. 26, respectively.
Figure 27: Accuracy vs. Magnification Factor across Methods with 2 models on Cifar10 dataset.
Figure 28: Accuracy vs. Magnification Factor across Methods with 3 models on Cifar10 dataset.
Figure 29: Accuracy vs. Magnification Factor across Methods with 5 models on Cifar10 dataset.
Figure 30: Accuracy vs. Magnification Factor across Methods with 7 models on Cifar10 dataset.
Figure 31: Accuracy vs. Magnification Factor across Methods with 10 models on Cifar10 dataset.
Figure 32: Uniform soup - Logits Ensemble Accuracy vs. Magnification Factor across different model numbers on Cifar10 dataset.
Similar to Cifar100, the 2 models, 3 models, 5 models, 7 models and 10 models accuracy vs. magnification factor on Cifar10 are shown via Fig. 27, Fig. 28, Fig. 29, Fig. 30 and Fig. 31, respectively. Moreover, we also visualize the uniform soup and logits ensemble accuracy vs. different magnification factors across different model numbers on the Cifar10 dataset in Fig. 32. We can observe similar phenomenons as discussed in the apepr. However, on the Cifar10, the gaps have different peak points than Cifar100, which is ×$\times 110$ (for Cifar100, it is ×$\times 100$).

### v4

@@ 6.2.1 Observations over Different Models
As shown in Fig. 3(a) for the merging of 10 models (the experiments for merging from 2 to 7 models are in the supplementary materials). For abbreviations, “Perf Ave” denotes the average performance of individual models; “deit_tiny” and “vit_tiny” denote deit_tiny_patch16_224 and vit_tiny_patch16_224 models, respectively. Different models show different merging/ensembling performances, but model ensemble on logits generally performs the best across different tested architectures He et al. (2016); Huang et al. (2017); Simonyan (2014); Steiner et al. (2021); Dosovitskiy (2020); Touvron et al. (2021). However, uniform soups perform the worst. We can also discover that greedy soups generally perform better than uniform soups due to their mechanism to greedily select the best performing model to merge. In fact, it guarantees that reasonably promising results can be achieved by only incorporating the first model (if the later-on merged model performance is not ideal). The greedy ensemble on both logits and features produces results that are comparable to greedy soups.
In general, adding more models into the merging process leads to better performance for all methods, including model-merging and ensembling. Interestingly, the ViT model, including 3 ViT-based architectures in the experiments, shows large gaps between uniformly averaging and logits ensembling, where such gaps suffer larger increases by adding more models. It may be caused by the architectural differences between convolutional neural networks and Vision Transformers (ViTs). CNNs are more adapted to the learning of local patterns within different scales of images, making CNN architectures ideal for merging. On the contrary, ViTs have global attention mechanisms, where the templates represent attention and correlations between tokens, making them sensitive for small changes. ViT models also show similar deterioration on other data (as shown by the DeiT on TinyImageNet and it can be found in the supplementary materials).
@@ 6.2.2 Observations over Different Datasets
Table 1: Performance Comparison for VGG19 with 10 Models Across Different Datasets. We adopt averaged AUROC ( Area Under Curve of Receiver Operating Characteristic) as the evaluation metric (normalized into 0 to 100 for better representation with other datasets) for the ChestXRay Wang et al. (2017); Ma et al. (2019) dataset; while accuracy (0% to 100%) for the rest of datasets.
@@
For the dataset-wise examination, we perform model-merging and ensembling on 7 prevalent datasets with different data sizes (TinyImageNet and ChestXRay are relatively large, and both with more than 100,000 images in the datasets). From Tab. 1, we can observe that the ensembling still generally performs better than merging. The models on different datasets have different performances due to the dataset characteristics. Similar to the model-wise comparison, with more models in the merging process (2 to 7 models merging results are in the supplementary materials), the better the performance for both merging and ensembling. The ensembling is more stable than merging in terms of the best and worst performance on different datasets. For example, on the PathMNIST Yang et al. (2021, 2023b), uniform soup does not work very well (32.42% accuracy, which is largely behind model ensemble methods) compared to its performance on the Cifar10/Cifar100 or Tiny ImageNet datasets. A similar situation shows on Celeba Liu et al. (2015): the uniformly average model only achieves 53.19%. Nevertheless, the greedy soup models can mitigate this problem (88.48% on PathMNIST and 93.06 on Celeba) as they can ensure a decent result by only considering the first model. These results may be caused by the contradictions between model weights.
@@ 6.3 Model Predictions and Weight Magnitudes
We further visualize the gaps between model-merging and ensembling in fig. 5(c). When more models are taken into account, the gaps become larger, but there is a peak point for the gap (×$\times 100$ is the peak for Cifar100). The experiments of merging/ensembling 2 to 10 models with different magnitude factors, as well as on different datasets and the gaps between merging and ensembling, are shown in the supplementary materials. A similar phenomenon is shown on different datasets, but with different peaks.
@@ 7 Conclusion
Appendix A More Figure Illustrations
Appendix B Experimental Settings
@@
For the experimental settings of model weight patterns. We train one single linear layer on both Cifar100 Krizhevsky (2009) and Tiny ImageNet Le and Yang (2015) datasets with momentum SGD as the optimizer. The initial learning rater is set to 0.01 and decreases every 20 epochs by 0.1. While visualization, the weights are wrapped back to the image size as described in the main paper (i.e. 100 $\times$ 3 $\times$ 32 $\times$ 32 for Cifar100).
In terms of averaging on weights versus averaging on features, we consider Uniform Soups and Greedy Soups Wortsman et al. (2022) for weight-averaged model-merging, which is the focus of our paper. We design the experiments of this section from two different angles: model-wise examination and dataset-wise examination of the model-merging and model ensemble performances. For the models considered, we have ResNet He et al. (2016) (ResNet18 to ResNet152), DenseNet Huang et al. (2017) (DenseNet121 to DenseNet201), VGG Simonyan (2014) (VGG11 to VGG19, including with and without batch normalization), ViT Steiner et al. (2021); Dosovitskiy (2020) and DeiT Touvron et al. (2021). For the dataset-wise examination, we conduct experiments on Cifar10/Cifar100, PathMNIST/DermaMNIST Yang et al. (2021, 2023b), Celeba Liu et al. (2015), Tiny ImageNet and ChestXRay Wang et al. (2017); Ma et al. (2019). The information of these datasets can be find in table 2. Specially, ChestXRay-14 is a chest X-ray imaging dataset designed for multi-label classification. It consists of 112,120 frontal chest X-ray images from 30,805 unique patients. The dataset includes annotations for 14 different lung disease categories. We use the official dataset split for training and evaluation. We demonstrate the average Area Under the Receiver Operating Characteristic Curve (AUROC) across 14 labels as the model performance for ChestXRay.
@@
Appendix C Theoretical Proofs of Weight Magnitude/Variance Expansion Effects Output Magnitude/Variance
@@
For the Property 1 and Theorem 1, according to Wang et al. (2024), we have:
@@
Appendix D More Visualization of Model Weight Patterns
D.0.1 Visualize Patterns Captured in Weights of 100 Classes for Cifar100 Dataset
@@
D.0.2 Visualization of Deep Convolutional Kernels Trained on Cifar100 Dataset
@@
D.0.3 Visualize Patterns Captured in Weights of 200 Classes for TinyImageNet Dataset
@@
Appendix E More Experimental Results of Averaging on Weights V.S. Averaging on Features
E.1 Merging Models Trained over Different Tasks
Datasets Data Size Model2 Uniform Soups Ens Lgt Ens Features
Cifar10 60,000 5.84 49.63 69.93 69.93
PathMNIST 107,180 1.29 24.25 69.88 70.00
DermaMNIST 10,015 5.77 60.37 69.06 69.36
Celeba 202,599 1.03 49.33 69.48 69.61
TinyImageNet 100,000 18.16 61.02 58.40 58.52
ChestXRay 112,120 15.18 61.38 68.18 68.33
We also explore model-merging between different well-trained models on different tasks, as shown in table 3. Model 1 is a VGG19 model trained on the Cifar100 with the classification accuracy of 70.37%, while Model 2 denotes the VGG19 model trained on different datasets, but evaluated on Cifar100 after merging/ensembling. Generally, uniform souping within Tiny ImageNet, where Model 2 was trained on, has the highest accuracy of 61.02% that is better than the ensemble on logits (58.40% accuracy). It may be because the Tiny ImageNet dataset is a general and relatively large dataset, forcing Model 2 to learn a diverse set of features. Interestingly, Cifar10 should be a similar one to Cifar100, but the souping result does not outperform the model ensembling. Maybe because it is too small. In contrast, the ChestXRay dataset should be very different from Cifar100, but it shows a reasonably decent improvement with souping compared to Model 2 along. It illustrates that the dataset size also matters. The souping performance on Celeba (Model 2 is trained on) is not good, may be caused by the Celeba dataset only containing human faces that are very different from Cifar100 (Model 1 is trained on).
E.2 2 Models, 3 Models, 5 Models and 7 Models for Model Merging/Ensembling Accuracy Comparison
Figure 17: (a) The 2 model merging/ensembling accuracy comparison between different architectures on Cifar100. (b) The 3 model merging/ensembling accuracy comparison between different architectures on Cifar100.
Figure 18: (a) The 5 model merging/ensembling accuracy comparison between different architectures on Cifar100. (b) The 7 model merging/ensembling accuracy comparison between different architectures on Cifar100.
The 2 models, 3 models, 5 models and 7 models for model merging/ensembling accuracy comparison between different architectures on Cifar100 are shown via fig. 17(a), fig. 17(b), fig. 18(a) and fig. 18(b), respectively. They all show that the models with ensembling on logits generally receive the best results and uniform soups receive the worst results. Besides, model-merging performs much worse results for the ViT models and the gaps between uniform soups and ensemble on logits are enlarged with more models considered in.
@@ E.1 Merging Models Trained over Different Tasks
Also, we show the dataset-wise exploration of 2 models, 3 models, 5 models and 7 models merging/ensembling in table 4, table 5, table 6 and table 7, respectively. The computation consumptions in Floating Point Operations (FLOPs) and Parameter Numbers are in table 8.
@@
We illustrate the ViT model (DeiT Tiny) on another dataset — Tiny ImageNet in table 9. We can also see the more models are merged, the worse results from uniform soups can get. Also, we can observe the greedy soups, greedy ensemble on logits and greedy ensemble on features are not changed. It demonstrates the model-merging between 2 or more ViT models will not contribute to the performance. Thus, it shows the ViT models are not suitable for model-merging in the tested scenarios.
Appendix F More Experimental Results of Model Magnitude Changes
F.1 Model-merging can mitigate weight magnitude issue on Tiny ImageNet
Moreover, we conduct the weight magnitude experiment on a relatively large dataset, Tiny ImageNet, in Tab. 10. From the table, we can observe that model-merging can still mitigate the weight magnifying problem to a certain extent. It can also outperform logit ensembling, and cannot outperform feature ensembling. But generally it cannot work very well as on Cifar100. It may be caused by the complexity of the Tiny ImageNet and as we discussed in the Sec. 5.3, the model-merging can work to some extent as it can act similarly as regularization. It can mitigate the weight magnitude increment issues, but with the cost of losing model expressiveness.
Table 10: Model-merging for mitigate weight magnitude issue with different VGG19 model numbers on Tiny ImageNet with the magnitude factor ×$\times 100$ (accuracy is the metric). This approach works for smaller data (Cifar10/Cifar100) as it has a regularization effect. However, it does not work for larger datasets such as Tiny ImageNet, because it sacrifices the expressiveness of the models.
2 models 59.87 5.60 6.70 6.99 1.59 6.94 6.45 1.52 6.96
3 models 59.85 5.44 7.15 7.05 0.78 7.38 6.78 0.78 7.21
5 models 59.89 5.07 6.68 6.83 0.30 6.95 6.46 0.35 7.21
7 models 59.90 4.80 6.77 6.80 0.24 6.73 6.26 0.21 6.74
10 models 59.90 4.81 6.86 6.92 0.13 6.92 6.22 0.11 6.95
F.2 2 Models, 3 Models, 5 Models and 7 Models Accuracy V.S. Different Magnification Factors
Figure 20: (a) Accuracy vs. Magnification Factor across Methods with 2 models on Cifar100 dataset. (b) Accuracy vs. Magnification Factor across Methods with 3 models on Cifar100 dataset.
Figure 21: (a) Accuracy vs. Magnification Factor across Methods with 5 models on Cifar100 dataset. (b) Accuracy vs. Magnification Factor across Methods with 7 models on Cifar100 dataset.
The 2 models, 3 models, 5 models and 7 models accuracy vs. different magnification factors on Cifar100 are shown via fig. 20(a), fig. 20(b), fig. 21(a) and fig. 21(b), respectively.
Figure 22: (a) Accuracy vs. Magnification Factor across Methods with 2 models on Cifar10 dataset. (b) Accuracy vs. Magnification Factor across Methods with 3 models on Cifar10 dataset.
Figure 23: (a) Accuracy vs. Magnification Factor across Methods with 5 models on Cifar10 dataset. (b) Accuracy vs. Magnification Factor across Methods with 7 models on Cifar10 dataset.
Figure 24: (a) Accuracy vs. Magnification Factor across Methods with 10 models on Cifar10 dataset. (b) Uniform soup - Logits Ensemble Accuracy vs. Magnification Factor across different model numbers on Cifar10 dataset.
Similar to Cifar100, the 2 models, 3 models, 5 models, 7 models and 10 models accuracy vs. magnification factor on Cifar10 are shown via fig. 22(a), fig. 22(b), fig. 23(a), fig. 23(b) and fig. 24(a), respectively. Moreover, we also visualize the uniform soup and logits ensemble accuracy vs. different magnification factors across different model numbers on the Cifar10 dataset in fig. 24(b). We can observe similar phenomenons as discussed in the apepr. However, on the Cifar10, the gaps have different peak points than Cifar100, which is ×$\times 110$ (for Cifar100, it is ×$\times 100$).